OpenStudy (anonymous):

What is the value of x? A. 1 B. 2 C. 3 D. 4

OpenStudy (anonymous):



OpenStudy (anonymous):

its an equilateral triangle, all sides will be equal hence 2x+2=4x-1 2x=3 x=1.5 i hope its correct

OpenStudy (anonymous):

@sriramkumar thats 2x+5. 2x+5 = 4x-1 6 = 2x and hence x = 3 ans is C that equality line looked like base of 2 for u

OpenStudy (anonymous):

exact answer gets a medal

OpenStudy (anonymous):

whatever @luckythebest I'm a teacher of process, not the solution :D anywasy thanks and sorry I was just mistaken :D